# A Novel Approach to Reduce Derivative Costs in Variational Quantum Algorithms

### v2

@@ V Second derivative and Hessian calculation
With a straightforward extension of Eq. (7), it can be shown that the second derivative, requires the evaluation of the cost function $f$ in four points Mari et al. (2021); Solinas et al. (2023). To implement this in the DM approach, we need four independent measurements of the observables $\hat{M}$. On the other hand, following the QNDM paradigm, we can just couple the system to the detector four times to store in the phase of the latest the same amount of information Solinas et al. (2023). Then, analogously to the first-order derivative, we perform a measure on the detector only once. For further details and the numerical simulations about the second derivatives with QNDM method, we address the interested reader to Appendix V. In this section, instead, we present the persistence of the advantages observed in first derivatives even when extending to second derivatives.
@@ V.1 Costs Analysis for second derivatives
In Table 2, we insert the costs for calculating second derivatives. We denote the resource cost function, $\mathcal{C}\left(g_{w,j}^{i}\right)$, for both the approaches ($i=QNDM,\,DM$) in terms of the number of gates required to compute the second derivative along the directions $w$ and $j$. These values are obtained from Refs. Solinas et al. (2023); Cerezo et al. (2021); Mari et al. (2021). As above, the costs depend on the number of Pauli strings $J$, the number of qubits $n$, the number of shots or repetitions $N_{i}$, and the number of logical operators needed to implement $U(\vec{\theta})$, denoted by $k$.
@@
DM $N_{DM}$ $\mathcal{C}\left(g_{w,j}^{DM}\right)=N_{DM}4J(k+n)$
QNDM $N_{QNDM}$ $\mathcal{C}\left(g_{w,j}^{QNDM}\right)=N_{QNDM}(7k+16Jn)$
@@
From Table 2, we observe that the cost of calculating the second derivative with DM is twice the cost for the first derivative with the same method. For QNDM, the variation in cost is due to the coefficient of $k$ (which increases from $3$ to $7$) and to the coefficient of $J$, which is twice the one for a single derivative.
@@
$\mathcal{C}\left(\frac{g_{w,j}^{DM}}{g_{w,j}^{QNDM}}\right)=\frac{4J}{7}\frac{N_{DM}}{N_{QNDM}}=\mathcal{O}(J^{2})$ (31)
where for the second derivative for DM approach, to reach the same precision obtained with the QNDM approach, the number of shots is proportional to $N_{DM}\propto N_{QNDM}J$. This is shown in Appendix A. The opposite regime is $k\ll nJ$. In this case, the resource ratio is given by
$\mathcal{C}\left(\frac{g_{w,j}^{DM}}{g_{w,j}^{QNDM}}\right)=\frac{k}{4n}\frac{N_{DM}}{N_{QNDM}}=\mathcal{O}(kJ)$ (32)
@@ VI Conclusions
This advantage can be attributed to two main features. First, as highlighted in Solinas et al. (2023), the use of a quantum detector to store information about the derivative of the cost function allows us to run the quantum circuit and measure only once. In contrast, the DM approach requires two circuits and two measurements (four circuits and four measurements for the Hessian). Second, introduced here for the first time, the statistical uncertainty (i.e., the mean square error) for the QNDM is independent of the number of Pauli strings $J$ appearing in the quantum observable, whereas for the DM approach, it increases linearly with $J$. Consequently, to achieve the same accuracy in the average values of the cost function for quantum observables composed of numerous Pauli strings $J$, the number of quantum circuit runs in the DM approach must increase.
The numerical simulations performed using the Qiskit framework have confirmed the analytical estimates. Even for small systems and simplified quantum circuits, we observe a considerable reduction in the resources needed to run the QNDM approach compared to the DM method. We anticipate that this advantage will further increase for more practical and complex problems, such as the simulation of chemical compounds, molecules, or small quantum systems Weber et al. (2014). These results position the QNDM approach as a valuable alternative for implementing Variational Quantum Algorithms (VQAs) on the next generation of noisy quantum computers.
@@ Acknowledgements.
PS and GM acknowledge financial support from INFN. This work was carried out while GM was enrolled in the Italian National Doctorate on Artificial Intelligence run by Sapienza University of Rome in collaboration with the Dept. of Informatics, Bioengineering, Robotics, and Systems Engineering, Polytechnic School of Genova.
@@ Appendix A Second derivatives
In this section, we present a detailed analysis for the second derivative. The parameter shift rule, Eq. (5), can be generalized to calculate the second derivative Mari et al. (2021).
$g_{w,j}$ $=$ $\frac{\partial^{2}f({\vec{\theta}})}{\partial\theta_{j}\partial\theta_{w}}=$
$=$ $\Big{[}f(\vec{\theta}+s(\hat{e}_{j}+\hat{e}_{w}))-f(\vec{\theta}+s(-\hat{e}_{j}+\hat{e}_{w}))$
-- $f(\vec{\theta}+s(\hat{e}_{j}-\hat{e}_{w}))+f(\vec{\theta}-s(\hat{e}_{j}+\hat{e}_{w}))\Big{]}\Big{[}2\sin^{2}s\Big{]}^{-1}.$
@@
$\frac{\partial^{2}f({\vec{\theta}})}{\partial\theta_{j}\partial\theta_{i}}$ $=$ $\Big{[}{\rm Tr}_{S}[U^{\dagger}(\vec{\theta}+s(e_{j}+e_{w}))\hat{M}U(\vec{\theta}+s(e_{j}+e_{w}))\rho_{S}^{0}-U^{\dagger}(\vec{\theta}+s(-e_{j}+e_{w}))\hat{M}U(\vec{\theta}+s(-e_{j}+e_{w}))\rho_{S}^{0}$
-- $U^{\dagger}(\vec{\theta}+s(e_{j}-e_{w}))\hat{M}U(\vec{\theta}+s(e_{j}-e_{w}))\rho_{S}^{0}+U^{\dagger}(\vec{\theta}-s(e_{j}+e_{w}))\hat{M}U(\vec{\theta}-s(e_{j}+e_{w}))\rho_{S}^{0}]\Big{]}\Big{[}4\sin^{2}s\Big{]}^{-1}.$
It follows that for DM method to calculate the values of $f(\vec{\theta}+s(\hat{e}_{j}+\hat{e}_{w}))$, we run the quantum circuit to implement the corresponding $U(\vec{\theta}+s(\hat{e}_{j}+\hat{e}_{w}))$ and then perform a projective measure for each Pauli string composing the observable $\hat{M}$ Mari et al. (2021). The same procedure is then implemented for $f(\vec{\theta}+s(-\hat{e}_{j}+\hat{e}_{w})),f(\vec{\theta}+s(\hat{e}_{j}-\hat{e}_{w})),f(\vec{\theta}-s(\hat{e}_{j}+\hat{e}_{w}))$, then the derivative can be easily extracted. In Fig. S1 the DMis represented schematically.
Figure S1: Quantum circuit of the implementation of the DM protocol. Here $U(\vec{\theta}+s\hat{e}_{j}+s^{\prime}\hat{e}_{w})$ is Eq. (1) with two parameters shifted in directions $\hat{e}_{j}$ and $\hat{e}_{w}$. The $J$ is equal to the number of Pauli strings and the values ±$\pm s$ and ±$\pm s^{\prime}$ are the shift for the parameter shift rule.
@@
$U(\vec{\theta}+s(e_{j}-e_{w}))$ $=$ $U_{1}$
$U^{\dagger}(\vec{\theta}+s(e_{j}-e_{w}))U(\vec{\theta}-s(e_{j}+e_{w}))$ $=$ $U_{2}$
$U^{\dagger}(\vec{\theta}-s(e_{j}+e_{w}))U(\vec{\theta}+s(-e_{j}+e_{w}))$ $=$ $U_{3}$
$U^{\dagger}(\vec{\theta}+s(-e_{j}+e_{w}))U(\vec{\theta}+s(e_{j}+e_{w}))$ $=$ $U_{4}.$ (S4)
Figure S2: Quantum circuit of the implementation of the QNDM protocol for second derivatives. Here, $H$ is the Hadamard gate, $S$ is the phase gate, $U_{1}=U(\vec{\theta}+s(e_{j}-e_{w}))$, $U_{2}=U^{\dagger}(\vec{\theta}+s(e_{j}-e_{w}))U(\vec{\theta}-s(e_{j}+e_{w}))$, $U_{3}=U^{\dagger}(\vec{\theta}-s(e_{j}+e_{w}))U(\vec{\theta}+s(-e_{j}+e_{w}))$, $U_{4}=U^{\dagger}(\vec{\theta}+s(-e_{j}+e_{w}))U(\vec{\theta}+s(e_{j}+e_{w}))$ and $U_{\pm}=\exp\{\pm i\lambda\hat{Z_{a}}\otimes\hat{M}\}$ is the system-detector coupling operator.
@@ A.1 Errors Analysis for Second Derivative
$\text{MSE}_{QNDM}(\hat{g}_{w,j})$ $=\frac{\lambda^{2}(\partial^{2}_{\lambda}\mathcal{G}_{\lambda})^{2}}{4}+\frac{\sigma_{D}^{2}}{16N_{QNDM}\sin^{4}s}\frac{1}{\lambda^{2}(1-(2P_{0}-1)^{2})}$ (S5)
$\text{MSE}_{DM}(\hat{g}_{w,j})$ $=\sum_{i}\frac{h^{2}_{i}\sigma_{s}^{2}}{4N_{DM}\sin^{4}s}.$ (S6)
Notice that the differences with the first derivative MSEs in Eqs. (27) and (23) are only in the denominator and for the second derivative, the MSEDM grows linearly with the number of Pauli strings $J$. Consequently, to achieve the same precision level as MSEQNDM in the DM approach, we need to increase the number of shots as a function of $J$. More specifically, we have that
$\frac{\text{MSE}_{DM}(\hat{g}_{w,j})}{\text{MSE}_{QNDM}(\hat{g}_{w,j})}=1\longrightarrow N_{DM}=N_{QNDM}\frac{4\sigma_{s}^{2}\lambda^{2}(1-(2P_{0}-1)^{2})}{\sigma^{2}_{D}}\sum_{i}h^{2}_{i}.$ (S7)
Where, similarly to Eq. (28), we keep the linearity in $J$, only with a modified multiplicative factor due to the differences in the denominators between the MSE for the first and the second derivatives.
@@ A.2 Cost Simulations for Second Derivative
In Section V.1, we found the asymptotic cost ratios $\mathcal{C}\left(g_{w,j}^{DM}/g_{w,j}^{QNDM}\right)$ for the two regimes: $k\ll nJ$ Eq. (32) and $k\gg nJ$ Eq. (31). In Figs. S4 and S3, we include simulations that illustrate how the theoretical results are also maintained for the second derivatives. For each value presented in the figures, we averaged over $L=50$ different realizations.
Figure S3: Regime $k\gg nJ$. $a)$ Number of resources $\mathcal{C}(g_{w,j}^{i})$ needed to estimate the second derivative of the cost function along the directions $j$ and $w$; here, $i=QNDM$ (orange line) and $i=DM$ (blue line). $\mathcal{C}(g_{w,j}^{i})$ is plotted in function of the number of logical operators $k$ for a fixed $nJ=300$. $b)$ Ratio between DM and QNDM resources numbers $\mathcal{C}\left(g_{w,j}^{DM}/g_{w,j}^{QNDM}\right)$ as a function of Pauli string $J$ each colored line corresponds at a fixed $k$. The quantum circuit is composed of $n=10$ qubits and the average is taken over $L=50$ realization as discussed in the main text. The number of shots for the QNDMis $N_{QNDM}=200$, for DM, it is calculated for each realization by (S7), such that the MSE errors are equal for both methods.
Panel a) in Fig. S3 shows the numbers of resources, $\mathcal{C}\left(g_{w,j}^{i}\right)$, in the function of the logical operators $k$, when we consider the limit of $k\gg nJ$. The simulations are done for $n=10$ qubits, fixed $nJ=300$, and fixed shots $N=200$. In the second panel b) we show the ratio of the resources as a function of the number of Pauli strings $J$ and for $k$ different numbers of logical operators in $U(\vec{\theta})$.
Figure S4: $a)$ The averaged number of resources $\mathcal{C}(g_{w,j}^{i})$ needed to estimate the second derivative of the cost function along the directions $j$ and $w$; here, $i=QNDM$ (orange line) and $i=DM$ (blue line). $\mathcal{C}(g_{w,j}^{i})$ is plotted in function of the number of Pauli strings $J$ for a fixed number of logical operators $k=0.5\cdot 10^{3}$. $b)$ Ratio between DM and QNDM resources numbers $\mathcal{C}\left(g_{w,j}^{DM}/g_{w,j}^{QNDM}\right)$ as a function of logical operators $k$ each coloured line corresponds at a fixed $nJ$. The quantum circuit is composed of $n=10$ qubits and the average is taken over $L=50$ realization as discussed in the main text. The number of shots for the QNDMis $N_{QNDM}=200$, for DM, it is calculated for each realization by (S7), such that the MSE errors are equal for both methods.
Panel a) in Fig. S4 shows the numbers of resources, $\mathcal{C}\left(g_{w,j}^{i}\right)$, in function of $nJ$. We are working in the limit $k\ll nJ$. The simulations are done for $n=10$ qubits, a fixed number of logical operators $k=0.5\cdot 10^{3}$, and fixed shots $N=200$. In the second panel b) we show the ratio of the resources as a function of logical operators $k$ for different values of the product $nJ$. The advantage of the QNDM over the DM approach is confirmed also for the second derivatives.

### v3

@@ V Second derivative and Hessian calculation
With a straightforward extension of Eq. (7), it can be shown that the second derivative, requires the evaluation of the cost function $f$ in four points Mari et al. (2021); Solinas et al. (2023). To implement this in the DM approach, we need four independent measurements of the observables $\hat{M}$. On the other hand, following the QNDM paradigm, we can just couple the system to the detector four times to store in the phase of the latest the same amount of information Solinas et al. (2023). Then, analogously to the first-order derivative, we perform a measure on the detector only once. For further details and the numerical simulations about the second derivatives with QNDM method, we address the interested reader to Appendix A. In this section, instead, we present the persistence of the advantages observed in first derivatives even when extending to second derivatives.
@@ V.1 Costs Analysis for second derivatives
In Table 2, we insert the costs for calculating second derivatives. We denote the resource cost function, $\mathcal{C}\left(g_{w,l}^{i}\right)$, for both the approaches ($i=QNDM,\,DM$) in terms of the number of gates required to compute the second derivative along the directions $w$ and $l$. These values are obtained from Refs. Solinas et al. (2023); Cerezo et al. (2021); Mari et al. (2021). As above, the costs depend on the number of Pauli strings $J$, the number of qubits $n$, the number of shots or repetitions $N_{i}$, and the number of logical operators needed to implement $U(\vec{\theta})$, denoted by $k$.
@@
DM $N_{DM}$ $\mathcal{C}\left(g_{w,l}^{DM}\right)=N_{DM}4J(k+n)$
QNDM $N_{QNDM}$ $\mathcal{C}\left(g_{w,l}^{QNDM}\right)=N_{QNDM}(7k+16Jn)$
@@
From Table 2, we observe that the cost of calculating the second derivative with DM is twice the cost for the first derivative with the same method. For QNDM, the variation in cost is due to the coefficient of $k$ (which increases from $3$ to $7$) and to the coefficient of $J$, which is twice the one for a single derivative. A detailed description of the implementation of both approaches for computing the second derivative is provided in Appendix A.
@@
$\mathcal{C}\left(\frac{g_{w,l}^{DM}}{g_{w,l}^{QNDM}}\right)=\frac{4J}{7}\frac{N_{DM}}{N_{QNDM}}=\mathcal{O}(J)$ (32)
The opposite regime is $k\ll nJ$. In this case, the resource ratio is given by
$\mathcal{C}\left(\frac{g_{w,l}^{DM}}{g_{w,l}^{QNDM}}\right)=\frac{k}{4n}\frac{N_{DM}}{N_{QNDM}}=\mathcal{O}(k)$ (33)
@@ VI Conclusions
This advantage, as highlighted in Solinas et al. (2023), arises from the use of a quantum detector to store information about the derivative of the cost function, allowing us to run the quantum circuit and perform measurements only once. In contrast, the DM approach requires executing two separate circuits, with each circuit needing to be repeated for the number of Pauli strings. For the second derivative case, this requirement increases to four circuits.
The numerical simulations performed using the $Qiskit$ framework have confirmed the analytical estimates. Even for small systems and simplified quantum circuits, we observe a considerable reduction in the resources needed to run the QNDM approach compared to the DM method. We anticipate that this advantage will further increase for more practical and complex problems, such as the simulation of chemical compounds, molecules, or small quantum systems Weber et al. (2014); Sawaya et al. (2024). These results position the QNDM approach as a valuable alternative for implementing Variational Quantum Algorithms (VQAs) on the next generation of noisy quantum computers.
@@ Acknowledgements.
PS, GM, and DM acknowledge financial support from INFN. This work was carried out while GM was enrolled in the Italian National Doctorate on Artificial Intelligence run by Sapienza University of Rome in collaboration with the Dept. of Informatics, Bioengineering, Robotics, and Systems Engineering, Polytechnic School of Genova.
@@ Appendix A Second derivatives
In this section, we present a detailed analysis of the second derivative. The parameter shift rule, Eq. (5), can be generalized to calculate the second derivative Mari et al. (2021).
$g_{w,l}$ $=$ $\frac{\partial^{2}f({\vec{\theta}})}{\partial\theta_{l}\partial\theta_{w}}=$
$=$ $\Big{[}f(\vec{\theta}+s(\hat{e}_{l}+\hat{e}_{w}))-f(\vec{\theta}+s(-\hat{e}_{l}+\hat{e}_{w}))$
-- $f(\vec{\theta}+s(\hat{e}_{l}-\hat{e}_{w}))+f(\vec{\theta}-s(\hat{e}_{l}+\hat{e}_{w}))\Big{]}\Big{[}2\sin^{2}s\Big{]}^{-1}.$
@@
$\frac{\partial^{2}f({\vec{\theta}})}{\partial\theta_{l}\partial\theta_{i}}$ $=$ $\Big{[}{\rm Tr}_{S}[U^{\dagger}(\vec{\theta}+s(e_{l}+e_{w}))\hat{M}U(\vec{\theta}+s(e_{l}+e_{w}))\rho_{S}^{0}-U^{\dagger}(\vec{\theta}+s(-e_{l}+e_{w}))\hat{M}U(\vec{\theta}+s(-e_{l}+e_{w}))\rho_{S}^{0}$
-- $U^{\dagger}(\vec{\theta}+s(e_{l}-e_{w}))\hat{M}U(\vec{\theta}+s(e_{l}-e_{w}))\rho_{S}^{0}+U^{\dagger}(\vec{\theta}-s(e_{l}+e_{w}))\hat{M}U(\vec{\theta}-s(e_{l}+e_{w}))\rho_{S}^{0}]\Big{]}\Big{[}4\sin^{2}s\Big{]}^{-1}.$
It follows that for DM method to calculate the values of $f(\vec{\theta}+s(\hat{e}_{l}+\hat{e}_{w}))$, we run the quantum circuit to implement the corresponding $U(\vec{\theta}+s(\hat{e}_{l}+\hat{e}_{w}))$ and then perform a projective measurement for each Pauli string composing the observable $\hat{M}$ Mari et al. (2021). The same procedure is then implemented for $f(\vec{\theta}+s(-\hat{e}_{l}+\hat{e}_{w})),f(\vec{\theta}+s(\hat{e}_{l}-\hat{e}_{w})),f(\vec{\theta}-s(\hat{e}_{l}+\hat{e}_{w}))$, then the derivative can be easily extracted. In Fig. S1 the DMis represented schematically.
Figure S1: Quantum circuit of the implementation of the DM protocol. Here $U(\vec{\theta}+s\hat{e}_{l}+s^{\prime}\hat{e}_{w})$ is Eq. (1) with two parameters shifted in directions $\hat{e}_{l}$ and $\hat{e}_{w}$. The $J$ is equal to the number of Pauli strings and the values ±$\pm s$ and ±$\pm s^{\prime}$ are the shift for the parameter shift rule.
@@
$U(\vec{\theta}+s(e_{l}-e_{w}))$ $=$ $U_{1}$
$U^{\dagger}(\vec{\theta}+s(e_{l}-e_{w}))U(\vec{\theta}-s(e_{l}+e_{w}))$ $=$ $U_{2}$
$U^{\dagger}(\vec{\theta}-s(e_{l}+e_{w}))U(\vec{\theta}+s(-e_{l}+e_{w}))$ $=$ $U_{3}$
$U^{\dagger}(\vec{\theta}+s(-e_{l}+e_{w}))U(\vec{\theta}+s(e_{l}+e_{w}))$ $=$ $U_{4}.$ (S4)
Figure S2: Quantum circuit of the implementation of the QNDM protocol for second derivatives. Here, $H$ is the Hadamard gate, $S$ is the phase gate, $U_{1}=U(\vec{\theta}+s(e_{l}-e_{w}))$, $U_{2}=U^{\dagger}(\vec{\theta}+s(e_{l}-e_{w}))U(\vec{\theta}-s(e_{l}+e_{w}))$, $U_{3}=U^{\dagger}(\vec{\theta}-s(e_{l}+e_{w}))U(\vec{\theta}+s(-e_{l}+e_{w}))$, $U_{4}=U^{\dagger}(\vec{\theta}+s(-e_{l}+e_{w}))U(\vec{\theta}+s(e_{l}+e_{w}))$ and $U_{\pm}=\exp\{\pm i\lambda\hat{Z_{a}}\otimes\hat{M}\}$ is the system-detector coupling operator.
@@ A.1 Errors Analysis for Second Derivative
$\text{MSE}_{QNDM}(\hat{g}_{w,l})$ $=\frac{\lambda^{2}(\partial^{2}_{\lambda}\mathcal{G}_{\lambda})^{2}}{4}+\frac{\sigma_{D}^{2}}{16N_{QNDM}\sin^{4}s}\frac{1}{\lambda^{2}(1-(2P_{0}-1)^{2})}$ (S5)
$\text{MSE}_{DM}(\hat{g}_{w,l})$ $=\sum^{J}_{i}=\frac{h^{2}_{i}\sigma_{s}^{2}}{4N_{DM}\sin^{4}s}.$ (S6)
@@ A.2 Cost Simulations for Second Derivative
In Section V.1, we found the asymptotic cost ratios $\mathcal{C}\left(g_{w,l}^{DM}/g_{w,l}^{QNDM}\right)$ for the two regimes: $k\ll nJ$ Eq. (33) and $k\gg nJ$ Eq. (32). In Figs. S4 and S3, we include simulations that illustrate how the theoretical results are also maintained for the second derivatives. For each value presented in the figures, we averaged over $L=50$ different realizations.
Figure S3: Regime $k\gg nJ$. $a)$ Number of resources $\mathcal{C}(g_{w,l}^{i})$ needed to estimate the second derivative of the cost function along the directions $l$ and $w$; here, $i=QNDM$ (orange line) and $i=DM$ (blue line). $\mathcal{C}(g_{w,l}^{i})$ is plotted in function of the number of logical operators $k$ for a fixed $nJ=240$. $b)$ Ratio between DM and QNDM resources numbers $\mathcal{C}\left(g_{w,l}^{DM}/g_{w,l}^{QNDM}\right)$ as a function of Pauli string $nJ$ each colored line corresponds at a fixed $k$. The quantum circuit is composed of $n=10$ qubits and the average is taken over $L=50$ realization as discussed in the main text. The number of shots for the QNDM is $N_{QNDM}=500$, for DM, it is calculated for each realization such that the MSE errors are equal for both methods.
Panel a) in Fig. S3 shows the numbers of resources, $\mathcal{C}\left(g_{w,l}^{i}\right)$, in the function of the logical operators $k$, when we consider the limit of $k\gg nJ$. The simulations are done for $n=10$ qubits, fixed $nJ=240$, and fixed shots $N=500$. In the second panel b) we show the ratio of the resources as a function of the number of Pauli strings $J$ and for $k$ different numbers of logical operators in $U(\vec{\theta})$.
Figure S4: $a)$ The averaged number of resources $\mathcal{C}(g_{w,l}^{i})$ needed to estimate the second derivative of the cost function along the directions $l$ and $w$; here, $i=QNDM$ (orange line) and $i=DM$ (blue line). $\mathcal{C}(g_{w,l}^{i})$ is plotted in function of the number of Pauli strings $J$ for a fixed number of logical operators $k=5\cdot 10^{3}$. $b)$ Ratio between DM and QNDM resources numbers $\mathcal{C}\left(g_{w,l}^{DM}/g_{w,l}^{QNDM}\right)$ as a function of logical operators $k$ each coloured line corresponds at a fixed $nJ$. The quantum circuit is composed of $n=10$ qubits and the average is taken over $L=50$ realization as discussed in the main text. The number of shots for the QNDMis $N_{QNDM}=500$, for DM, it is calculated for each realization such that the MSE errors are equal for both methods.
Panel a) in Fig. S4 shows the numbers of resources, $\mathcal{C}\left(g_{w,l}^{i}\right)$, in function of $nJ$. We are working in the limit $k\ll nJ$. The simulations are done for $n=10$ qubits, a fixed number of logical operators $k=5\cdot 10^{2}$, and fixed shots $N=500$. In the second panel b) we show the ratio of the resources as a function of logical operators $k$ for different values of the product $nJ$. The advantage of the QNDM over the DM approach is confirmed also for the second derivatives.